# GL1-02-06-24-2-supplement

### v2

@@ Abstract
In this paper, we develop a Difference-in-Differences model for discrete, ordered outcomes, building upon elements from a continuous Changes-in-Changes model. We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion, substance abuse, or domestic violence, where too many “false zeros”, or more broadly, underreporting are likely. We provide characterizations for distributional parallel trends, a concept central to our approach, within a general threshold-crossing model framework. In cases where outcomes are assumed to be reported correctly, we propose a framework for identifying and estimating treatment effects across the entire distribution. This framework is then extended to modeling underreported outcomes, allowing the reporting decision to depend on treatment status. A simulation study documents the finite sample performance of the estimators. Applying our methodology, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The results indicate small, but significant increases in consumption probabilities at each level. These effects are further amplified upon accounting for misreporting.
@@ 1 Introduction
The Difference-in-differences (DiD) method is widely recognized in the econometric literature as a quasi-experimental technique for identifying causal mean parameters. While this method has been central in understanding the average effects of reforms or policy changes, there has been a growing interest in capturing the full spectrum of counterfactual outcomes as policy interventions may have multifaceted impacts across the entire distribution. Consequently, quasi-experimental methods exploiting pre- and post-treatment data have emerged over the past two decades to identify and estimate such distributional treatment effects (see e.g., Roth and Sant’Anna, 2023, for an overview). A prominent example of this more comprehensive approach is the Changes-in-Changes (CiC) model developed by Athey and Imbens (2006). This model is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation, and has gained popularity in recent applied (e.g., Gregory and Zierahn, 2022; Assunçao et al., 2023) and theoretical (e.g., Huber et al., 2022; Ghanem et al., 2023) literature. The key assumption is that the distribution of unobservables governing counterfactual outcomes remains stable over time within the treated and comparison groups.
@@
This paper presents a generalization of the the commonly used Logit or Probit DiD method to discrete, ordered outcomes, building upon elements from a continuous CiC model. Central to our approach is the concept of distributional parallel trends, which we leverage for the latent outcome variable within a general threshold-crossing model class, using it as our key identification assumption. Our modeling approach is distinct from and nonnested within the discrete CiC model of Athey and Imbens (2006), even though it does share the features of the continuous CiC approach. As a second generalization, we consider discrete outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case. This phenomenon of underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited. Common examples include the individual consumption of “social bads” such as marijuana (cannabis) and alcohol, and sensitive issues like abortion or domestic violence incidents. These behaviors or events often lead to a significant proportion of “false zeros” in survey data, due to individuals’ reluctance to report such incidents truthfully. [^1] For example, Brown et al. (2018) found that the actual incidence of marijuana consumption in an Australian survey was nearly double the reported figure (23% vs. 12.2%), while Caudill and Mixon (2005) identified about 20 percentage points underreporting in cheating behavior among undergraduate students in a U.S. survey. See also Nguimkeu et al. (2019) for other examples. Such partial observability, without additional constraints on the Data Generating Process, poses significant challenges for both (point) identification and estimation.
The paper makes several contributions to the literature. Firstly, for the case where outcomes are assumed to be reported correctly, we propose a framework for identification and estimation of their causal effects across the entire distribution. Specifically, we consider a general ordered choice model with strictly stationary error terms, and demonstrate that the distributional parallel trends assumption is a natural choice for discrete outcomes. We provide two characterizations of distributional parallel trends on the latent outcomes within this threshold-crossing framework: we establish that distributional parallel trends hold if and only if a “parallel trends”-type assumption holds for the functional median (mean) and the scaling parameters. The latter holds if and only if these two functional parameters admit a distinct form of separability in group and time. These characterizations are of independent interest and provide a direct pathway to two key point identification results of the paper. In particular, we show that distributional parallel trends point identifies the counterfactual distribution of the latent outcome in a parametric ordered choice setup when no other covariates are present. We also demonstrate that, when additional covariates are present, the counterfactual distribution are point identified even in the absence of distributional assumptions on the error term, i.e., when the model is semiparametric. Our results nest the commonly used Logit or Probit DiD for discrete responses (see Puhani, 2012; Wooldridge, 2023), which feature in the applied literature either as main specification for binary (Meyer and Rosenbaum, 2001; Finkelstein, 2007; Mayer et al., 2014) and ordered choice outcomes (e.g., Sequeira, 2016), or as “robustness” specification (e.g., Marie and Zölitz, 2017; Hvide and Jones, 2018; Hoekstra and Sloan, 2022). These models typically draw on the conventional parallel trends that exclusively affect the mean parameters. Moreover, when additional covariates are present, our identification result can directly be translated into an estimation procedure that accommodates the discrete nature of the outcome. This is in contrast to Athey and Imbens (2006), where the proposed estimators are not directly tailored to these types of outcomes and do not allow for structural estimation.
@@
As a second contribution, this paper extends the DiD framework to account for underreporting, a scenario leading to partially observable data, where the true outcomes are correctly recorded if and only if they are both actual and reported (see Poirier, 1980). We generalize the conventional ”false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only. We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting the decisions, nor on their evolution across groups and over time. In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000a). Addressing the identification and estimation challenges posed by such partially observed outcome models, we propose a set of sufficient conditions for the identification of causal parameters. Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017; Han and Lee, 2019), is notably more complex in its analysis due to the partial observability. We demonstrate that having an exclusion restriction in each equation (outcome and reporting), where the excluded variables are continuous, along with the real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated. This real analytic assumption, a strong form of smoothness imposed on the copula and the marginal distribution functions of the unobservables, has been used for identification in various contexts including quantile sample selection (Arellano and Bonhomme, 2017), triangular models with endogenous binary treatment (Pereda-Fernández, 2023), and random coefficient discrete choice models (Fox et al., 2012; Fox and Gandhi, 2016). While not being “constructive” and lacking a direct pathway for estimation, it permits the use of a wide range of well-established marginal distribution and copula functions, thereby offering practical flexibility. On the other hand, two exclusion restrictions are required to identify the parameters in both outcome and reporting equations. This is common in multi-index outcome models (e.g., Donkers and Schafgans, 2008; Lewbel and Lin, 2022), but may be hard to find in some applications. However, we note that this requirement for two exclusion restrictions can be relaxed if the focus is solely on the parameters of one of these equations. In addition, we discuss possibilities to device specification tests that allow to check the functional specification in practice. Crucially, our approach accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (e.g., Arulampalam et al., 2017). This aspect is particularly pertinent in our empirical application, where we examine how reporting behavior may shift in response to treatment.
The theoretical identification results are complemented by an extensive Monte Carlo study that investigates the finite sample performance across various exclusion restriction scenarios and distributional choices related to the real analytic copula function. Corroborating that parameters of the partially observable outcome model remain unidentified without exclusion restrictions as postulated by Poirier (1980), our study reveals a relatively sizable and slow-decreasing bias in the Maximum Likelihood (ML) estimator for certain parameters, when a single continuous excluded variable is employed in only one of the equations. This pattern persists across both Frank and Clayton copula functions. Crucially, the inclusion of two excluded variables, one per equation, substantially improves finite sample performance. This leads to a rapid decline in the bias of the ML estimator for all model parameters, offering practical guidance for applying these methods in real-world scenarios.
@@
Finally, as an empirical contribution, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The focus on the last group is particularly relevant as they are widely believed to be most vulnerable to the associated health risks of marijuana consumption. Unlike existing studies which examined the effects of legalization of medicinal or recreational marijuana usage on the mean (cf. Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we consider the effects of the legalization of recreational marijuana use on the entire consumption distribution of the aforementioned high-school population in the treated states of California, Maine, Nevada, and Massachusetts, thus allowing for the possibility of heterogeneous effects of the reforms. Moreover, in contrast to the existing literature, we address the concerns that consumption behavior may not always be reported truthfully, and that reporting behavior may have been affected by legalization through shifts in stigma, public perception, or general attitudes towards marijuana consumption. Our analysis reveals that, while initial findings without covariates indicate no significant changes, incorporating covariates uncovers a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point at each level. Notably, after accounting for misreporting, the effect sizes increase by more than 50% at each level. Despite these shifts in consumption, our findings do not corroborate significant changes in reporting behavior.
The paper is organized as follows. Section 2 outlines the model setup, detailing the observation rule for self-reported outcomes and defining the distributional treatment effects of interest. Section 3 introduces the concept of distributional parallel trends, providing necessary and sufficient conditions for them to hold within a general threshold-crossing framework. Section 4 addresses point identification of model parameters when outcomes are correctly reported, starting with the scenario without covariates and then moving to a semiparametric model with unspecified error term distributions. Section 5 extends the framework to account for underreporting, establishing point identification in a corresponding semiparametric model. Section 6 is dedicated to point identification of distributional treatment effects on the treated, both with and without misreporting. Section 7 discusses both parametric and semiparametric estimation for these distributional treatment effects. Section 8 presents extensive Monte Carlo simulations to assess the finite sample behavior of the ML estimator in the scenario of underreported outcomes, covering various exclusion restriction scenarios and contrasting different index and error variances. Section 9 empirically investigates the effects of recreational marijuana legalization on the consumption behavior of 8th grade high-school students. Section 10 concludes. Finally, an alternative characterization of distributional parallel trends can be found in Appendix A, while Appendix B contains the proofs of all propositions in the paper. A proof for the result of Appendix A, details on the calculation of the standard errors, an alternative estimation procedure as well as additional simulation and empirical results can be found in the supplementary material.
@@ 2 Setup
Let $Y(1)$ and $Y(0)$ denote the potential outcomes for individuals that take on a finite number of $J+1$ values (in a pre-specified period of time) with and without treatment, respectively. In light of the empirical application of Section 9, we will refer to $Y(1)$ and $Y(0)$ as potential consumption levels in the following. Also, for simplicity, we will focus on the case of two groups and two time periods exclusively. The extension to several (pre-treatment) periods is immediate at the cost of more cumbersome notation. We define $G\in\{0,1\}$ as the group membership indicator (where group 1 is the treated group) and $T\in\{0,1\}$ as the time period indicator (where period 1 is after the treatment). The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if $(G,T)=(1,1)$, and $D=0$ otherwise. To streamline the notation, we adopt the shorthand as in Athey and Imbens (2006):
@@
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”. [^2] We will adopt the same notation for all other random variables in this paper.
As mentioned in the introduction, a prevalent issue in analyzing sensitive behaviors, such as the consumption of “social bads,” is the presence of “false zeros” or, more broadly, underreporting. Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019). This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf. Poirier, 1980). To formalize this idea, let $R(d)$ denote the potential “reporting intentions” for $d=0,1$. Importantly, note that this formulation allows reporting behavior to change in response to the treatment status $d$. For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully. Using similar shorthand as before, we continue to denote
$R_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}R(d)\mid G=g,T=t$
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$. Here, we interpret $R_{gt}(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully. The potential outcome of interest, $C_{gt}(d)$, $d=0,1$, is then given by
$C_{gt}(d)=\min\left\{Y_{gt}(d),R_{gt}(d)\right\}.$ (1)
Thus, respondents will truthfully report their consumption $Y_{gt}(d)$ if and only if it is less than or equal to the reporting intention $R_{gt}(d)$. For instance, if $Y_{gt}(0)=2$ and $R_{gt}(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2. It is worth noting that this formulation of underreporting behavior in (1) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019). Building on this framework and (1), we may also define the actual observed outcome $C_{gt}$ in the standard manner as a linear combination of the potential outcomes $C_{gt}(0)$ and $C_{gt}(1)$ as:
$C_{gt}=C_{gt}(1)D+C_{gt}(0)(1-D),$
where $C_{gt}(d)=\min\{Y_{gt}(d),R_{gt}(d)\}$ for $d\in\{0,1\}$. The above definition bears similarities with the binary selection setup in Klein et al. (2015). However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)). That is, unlike a selection setup where the selection indicator is fully observed, $R_{gt}(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions. More specifically, both the actual consumption and reporting decisions are observable if and only if individuals have reported the highest consumption level.
Throughout the paper, our primary interest lies in identifying the distributional treatment effects on the treated (DTTs), focusing primarily on the impact of marijuana legalization on the drug consumption of adolescents residing in treated states. That is, in the presence of additional covariates $\mathbf{X}_{gt}$ (recall that the subscripts highlight the possible dependence on the group membership $G=g$ and time period $T=t$), we aim to identify:
$\tau_{11}^{c}(j\mid\mathbf{x}):=$ $\Pr\left(Y_{11}(1)=j\mid\mathbf{X}_{11}=\mathbf{x}\right)-\Pr\left(Y_{11}(0)=j\mid\mathbf{X}_{11}=\mathbf{x}\right),$ (2)
for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}_{11}$, i.e. values from the covariate distribution in the treated group after the treatment. Alternatively, we may also be interested in the unconditional distributional treatment effect on the treated. This can, depending on whether variation in $\mathbf{X}_{gt}$ plays essential roles in the identification arguments (see Section 4), either be directly addressed with an unconditional version of (2) or be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
[…]
where $\mathcal{X}_{gt}$ denotes the support of $\mathbf{X}_{gt}$ and ${F}_{\mathbf{X}_{gt}}(\cdot)$ the covariate distribution function for a given group $g$ and time period $t$. Another relevant policy parameter of interest is the distributional treatment effect for the treated on the reporting intention:
$\tau^{r}_{11}(j\mid\mathbf{z}):=$ $\Pr\left(R_{11}(1)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right)-\Pr\left(R_{11}(0)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right),$ (3)
for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}_{11}$, where $\mathbf{Z}_{11}$ denotes a covariate vector nonnested within $\mathbf{X}_{11}$, see Section 5. The parameter in (3) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment. This may also shed light on how factors like stigma and shame influence the reporting behavior.
Regardless of the estimand of interest, note that the first quantity on the RHS of (2) is generally nonparametrically identified from the data when $Y_{11}$ is observable, whereas the second probability is a counterfactual distribution that cannot be identified without further assumptions. As is customary in the DiD or the synthetic control literature, we therefore focus on modeling the counterfactual outcome $Y_{11}(0)$. To do so, for each group and time period $(g,t)\in\{0,1\}$, we consider a general threshold-crossing model framework for the counterfactual outcomes and introduce a latent outcome variable $Y^{*}_{gt}(0)$:
$Y_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}_{gt}(0)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}_{gt}(0)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}_{gt}(0)\leq\kappa_{3},\end{array}\right.$
[…]
$R_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}_{gt}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}_{gt}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}_{gt}(0)\leq\iota_{3},\end{array}\right.$ (7)
3 Characterizing Distributional Parallel Trends
@@
We start by considering the setup in the absence of misreporting, i.e. $C_{gt}(d)=Y_{gt}(d)$ for $d\in\{0,1\}$ almost surely, and provide characterizations of the distributional parallel trends in a general threshold-crossing setup where the unobservable distribution is invariant over time within each group. Specifically, recall that $\mathbf{X}_{gt}$ denotes a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively. Then, letting +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\mid x>0\}$, we define the latent consumption $Y_{gt}^{\ast}(0)$ to take the following form:
@@
where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is strictly stationary within each group given $\mathbf{X}_{gt}$ with zero median (or mean) and unit variance, and has an unknown, continuous, and strictly monotonic conditional cumulative distribution function (CDF) $F_{\varepsilon_{gt}\mid\mathbf{X}_{gt}}$.
Here, $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional median (mean) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution. [^4] To clarify the notation, the functions $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ are assumed to be identical to all units within the group $g$ and in period $t$, while the random variables $Y^{*}_{i,gt}(0)$, $\mathbf{X}_{i,gt}$, $\varepsilon_{i,gt}$ can vary across units. We drop the subscript $i$ to ease the notation. Note that $\eta_{gt}$ and $\lambda_{gt}$ can also be treated as trivial functions if one is interested in unconditional distributions. While the representation of $Y^{*}_{gt}(0)$ in (8) is w.l.o.g. for any given $g\in\{0,1\}$ and $t\in\{0,1\}$ (provided minor regularity conditions on the moments), the strict stationarity on the other hand postulates that the conditional distribution of $\varepsilon_{gt}$ must not change over time within each group. The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of both the DiD and the CiC approaches. Finally, note that Assumption 1 also nests various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) as well as nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
We next outline the definition of (conditional) distributional parallel trends for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper:

### v3

# GL1-22-11-24-2-supplement

@@ Abstract
In this paper, we develop a Difference-in-Differences model for discrete, ordered outcomes, building upon elements from a continuous Changes-in-Changes model. We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion or substance abuse, where too many “false zeros”, or more broadly, underreporting are likely. We start by providing a characterization for parallel trends within a general threshold-crossing model. We then propose a partial and point identification framework for different distributional treatment effects when the outcome is subject to underreporting. Applying our methodology, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The results indicate small, but significant increases in consumption probabilities at each level. These effects are further amplified upon accounting for misreporting.
@@ 1 Introduction
The Difference-in-differences (DiD) method is widely recognized in the econometric literature as a quasi-experimental technique for identifying causal mean parameters. While this method has been central in understanding the average effects of reforms or policy changes, there has been a growing interest in capturing the full spectrum of counterfactual outcomes as policy interventions may have multifaceted impacts across the entire distribution. Consequently, quasi-experimental methods exploiting pre- and post-treatment data have emerged over the past two decades to identify and estimate such distributional treatment effects (see e.g., Roth and Sant’Anna, 2023, for an overview). A prominent example of this more comprehensive approach is the Changes-in-Changes (CiC) model developed by Athey and Imbens (2006).
@@
This paper presents a CiC model for discrete, ordered outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case. Underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited, posing significant challenges for both model identification as well as estimation and inference. Common examples include the individual consumption of “social bads” such as marijuana (cannabis) and alcohol, and sensitive issues like abortion or domestic violence incidents. [^1] For example, Brown et al. (2018) found that the actual incidence of marijuana consumption in an Australian survey was nearly double the reported figure (23% vs. 12.2%), while Caudill and Mixon (2005) identified about 20 percentage points underreporting in cheating behavior among undergraduate students in a U.S. survey. See also Nguimkeu et al. (2019) for other examples.
The paper makes several contributions to the literature. Firstly, we consider a general ordered choice model with underreporting and with strictly stationary error terms along the lines of the continuous CiC model, and provide a characterization of parallel trends on the latent outcomes within a threshold-crossing model framework: we establish that this type of “Changes-in-Changes” setup holds if and only if a parallel trends-type assumption holds for the functional mean (median) and the scaling parameters. In the binary case and under distributional assumptions, this characterization can nest the commonly used Logit or Probit DiD for binary responses (see Puhani, 2012; Wooldridge, 2023) as these models typically draw on the conventional parallel trends that exclusively affect the mean parameters.
As a second contribution, we derive sharp, nonparametric bounds for the (counter)factual outcome distribution of the treatment group after treatment in the presence of underreporting. We show that, without further restrictions, the bounds on the counterfactual distribution remain uninformative even if an instrumental exclusion restriction for the consumption is available. By contrast, leveraging the CiC structure alongside with an instrumental exclusion restriction and an upper bound on the cumulative misreporting probability, we demonstrate that more informative, tight bounds can potentially be achieved. Our bounds extend the partial identification results of Athey and Imbens (2006) for the discrete outcome case to underreporting. Moreover, the results also complement recent results of Mondal and Wang (2024) for binary outcome models under misreporting by adapting their findings to an ordered choice setup. A suitable estimation and inference procedure for the nonparametric bounds is discussed in the supplementary material.
@@
As a third contribution, we complement the above findings with a set of point identification results for a more structural approach that explicitly models the consumption and reporting behavior, respectively. Specifically, this model generalizes the conventional “false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only. We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting decisions, nor on their evolution across groups and over time. In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000). Our approach also accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (Arulampalam et al., 2017). We address the identification and estimation challenges posed by the partially observed outcome model by proposing a set of sufficient conditions for the identification of differnt causal parameters of interest for consumption and reporting behavior. Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017), is notably more complex in its analysis due to the partial observability. We demonstrate that having an exclusion restriction in either the reporting or the consumption equation that exhibits some continuous variation, along with a real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated for consumption or the reporting behavior, respectively. This real analytic assumption, a strong form of smoothness imposed on the copula and the marginal distribution functions of the unobservables, has been used for identification in various contexts including quantile sample selection (Arellano and Bonhomme, 2017), triangular models with endogenous binary treatment (Pereda-Fernández, 2023), and random coefficient discrete choice models (Fox et al., 2012; Fox and Gandhi, 2016).
@@
Finally, as an empirical application, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The focus on the last group is particularly relevant as they are widely believed to be most vulnerable to the associated health risks of marijuana consumption. Unlike existing studies which examined the effects of legalization of medicinal or recreational marijuana usage on the mean (cf. Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we consider the effects of the legalization of recreational marijuana use on the entire consumption distribution of the aforementioned high-school population in the treated states of California, Maine, Nevada, and Massachusetts, thus allowing for the possibility of heterogeneous effects of the reforms. Moreover, in contrast to the existing literature, we address the concerns that consumption behavior may not always be reported truthfully, and that reporting behavior may have been affected by legalization through shifts in stigma, public perception, or general attitudes towards marijuana consumption. Our analysis reveals that, while initial findings without covariates indicate no significant changes, incorporating covariates uncovers a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point at each level. Notably, after accounting for misreporting, the effect sizes increase by more than 50% at each level. Despite these shifts in consumption, our findings do not corroborate significant changes in reporting behavior.
The paper is organized as follows. Section 2 outlines the model setup. Section 3 introduces the concept of distributional parallel trends, providing necessary and sufficient conditions for them to hold within a general threshold-crossing framework. Section 4 extends the framework to account for underreporting, establishing our main identification results. Section 5 empirically investigates the effects of recreational marijuana legalization on the consumption behavior of 8th grade high-school students. Section 6 concludes.
The supplementary material contains auxiliary results for the partial identification model (Section S.1), the proofs of all formal results (Section S.2), an outline of estimation and inference for the nonparametric bounds as well as for the semiparametric approach modeling the reporting behavior explicitly (Section S.3), a Monte Carlo study where we examine the finite sample behavior of the estimators under different sets of exclusion restrictions (Section S.4), additional empirical results (Section S.5). Finally, the supplement also provides a different Copula decomposition as a basis to construct different treatment effects (Section S.6).
@@ 2 Setup
In order to introduce the CiC framework for the ordered choice model and the parallel trends characterization in the next section, we start by considering the setup in the absence of underreporting. Let $Y(1)$ and $Y(0)$ denote the potential outcomes for individuals that take on a finite number of $J+1$ values (in a pre-specified period of time) with and without treatment, respectively. In light of the empirical application of Section 5, we will refer to $Y(1)$ and $Y(0)$ as “true” potential consumption levels in the following. For simplicity, we will focus on the case of two groups and two time periods exclusively. The extension to several (pre-treatment) periods is immediate at the cost of more cumbersome notation. We define $G\in\{0,1\}$ as the group membership indicator (where group 1 is the treated group) and $T\in\{0,1\}$ as the time period indicator (where period 1 is after the treatment). The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if and only if $(G,T)=(1,1)$, and $D=0$ otherwise.
Our primary goal lies in identifying distributional treatment effects on the treated (DTTs) for the outcome of interest, such as the effects of marijuana legalization on different levels of drug consumption among adolescents residing in treated states. That is, in the presence of additional covariates $\mathbf{X}$, we aim to identify:
[…]
for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}$ (conditional on $G=1$ and $T=1$), i.e., values from the covariate distribution in the treated group after the treatment. To distinguish the current focus from alternative treatment effects on reporting which will be introduced in Section 4, we use the superscript “c” to indicate “consumption”. Furthermore, we may also be interested in the unconditional distributional treatment effect on the treated. This can be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
[…]
where $\mathcal{X}_{11}$ denotes the support of $\mathbf{X}$ given $G=1$ and $T=1$, and ${F}_{\mathbf{X}|D=1}(\cdot)$ the covariate distribution function for the treated group after treatment.
In the absence of underreporting, the first quantity on the right-hand side RHS of (1) is generally nonparametrically identified from the data, while the second term represents a counterfactual distribution that cannot be identified without further assumptions. However, this condition is rarely satisfied when survey data on sensitive behaviors are involved, as such data are often subject to underreporting. Consequently, the presence of underreporting further complicates the identification of both quantities.
Given our primary interest is modeling self-reported ordered outcomes, we adopt a general threshold-crossing model framework for the counterfactual outcomes by introducing a latent outcome variable $Y^{*}(d)$, $d\in\{0,1\}$:
$Y(d)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}(d)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}(d)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}(d)\leq\kappa_{3},\end{array}\right.$
[…]
3 Parallel Trends for Ordered Outcomes
As is customary in the DiD or the synthetic control literature, we start by modeling the potential outcomes $Y(0)$. To streamline the notation, we adopt the shorthand from Athey and Imbens (2006) in what follows and write:
@@ 3 Parallel Trends for Ordered Outcomes
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”. We will adopt the same notation for all other random variables as well. Thus, let $\mathbf{X}_{gt}$ denote a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively. Then, letting +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\mid x>0\}$, we define the latent consumption $Y_{gt}^{\ast}(0)$ to take the following form:
@@
where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is continuously distributed and strictly stationary within each group given $\mathbf{X}_{gt}$.
Here, when one w.l.o.g. normalizes the mean (median) and variance of $\varepsilon_{gt}$ to zero and one, then $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional mean (median) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution. [^4] To clarify the notation, the functions $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ are assumed to be identical to all units within the group $g$ and in period $t$, while the random variables $Y^{*}_{i,gt}(0)$, $\mathbf{X}_{i,gt}$, $\varepsilon_{i,gt}$ can vary across units. We drop the subscript $i$ to ease the notation. Note that $\eta_{gt}$ and $\lambda_{gt}$ can also be treated as trivial functions if one is interested in unconditional distributions. While the representation of $Y^{*}_{gt}(0)$ in (2) is w.l.o.g. for any given $g\in\{0,1\}$ and $t\in\{0,1\}$ (provided minor regularity conditions on the moments), the strict stationarity on the other hand postulates that the conditional distribution of $\varepsilon_{gt}$ must not change over time within each group. The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of the CiC approach, which is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation (e.g., Assunçao et al., 2023; Huber et al., 2022). Finally, note that Assumption 1 encompasses various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) and is closely related to nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
We next outline the definition of a CiC model for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper. For a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, define its “generalized inverse” by
$F^{-1}:[0,1]\rightarrow\mathbb{R}\cup\left\{-\infty,\infty\right\},\ q\mapsto F^{-1}\left(q\right)$ $:=\inf\left\{y\in\mathcal{Y}\mid q\leq F\left(y\right)\right\},$
where we adopt the convention $\inf\{\emptyset\}=\infty$ and $\inf\{\mathbb{R}\}=-\infty$.